Aneku has 21 coins whose total value is Sh. 72, there are twice as many five shillings coins as there are 10 shilling coins. The rest are one shilling coins. Find the number of ten shilling coins that Aneku has
Answer by stanbon(75887)

Aneku has 21 coins whose total value is Sh. 72, there are twice as many five shillings coins as there are 10 shilling coins. The rest are one shilling coins. Find the number of ten shilling coins that Aneku has
---------------------
Quantity Eq.: f + t + n = 21
Value Eq:::::5f +10t +n = 72
Quantity Eq.: f = 2t
---------------------------------
Rearrange:
f + t + n = 21
5f +10t +n = 72
f - 2t +0 = 0
----------------------
I used a matrix method to get:
# of 5 schilling = 6
# of 10 schilling= 3
# of 1 schilling = 12
==============================
Cheers,
Stan H.
Reply to stanbon@comcast.net